OpenStudy (anonymous):

A crane lifts a 425 kg steel beam vertically a distance of 66 m. How much work does the crane do on the beam if the beam accelerates upward at 1.8 m/s2? Neglect frictional forces

OpenStudy (alfie):

How's work defined?

OpenStudy (anonymous):

F=ma = 425kg x 1.8m/s² =blablabla work done = force times distance=F x 66m=bla bla bla...

OpenStudy (king):

W=Fd=mad=425*1.8*66 =50490 J

OpenStudy (anonymous):

not correct